OpenStudy (anonymous):

Which of the following is the vertex of the function? y=x^2-6x+9 (3, 0) (-3, 36) (4, 1) (1, 4)

OpenStudy (anonymous):

first coordinate is \[-\frac{b}{2a}=-\frac{-6}{2\times 1}=3\]

OpenStudy (anonymous):

do u have to turn it into vertex form?